OpenStudy (anonymous):

Assume that y varies inversely as x. If y = -6 when x = -2, find y when x = 5. A. y = 1.7 B. y = 2.4 C. y = 3

OpenStudy (anonymous):

is that how its phrased exactly?

OpenStudy (anonymous):

Assume that y varies inversely as x?

OpenStudy (anonymous):

Yes and I dont get it

OpenStudy (merchandize):

yalpha1/x y=k/x -6=x/-2 k=12 x=5 , k=12 y=k/x y=12/5

OpenStudy (merchandize):

y=2.4